OpenStudy (anonymous):

are these answers correct 1) {-15}+{23}= >>>8 yes or no if correct 2) [-78]-[-12]=>>>66 yes or no 3) [17][-2]=>>>8.5 yes or no 4) [-45]/[9] >>-5 yes or no

OpenStudy (anonymous):

1) Yes 2) No 3) No 4) Yes

OpenStudy (anonymous):

well if 2 and 3 are wrong what is it

OpenStudy (anonymous):

2) -78-(-12) = -78+12=-66 3) (17)(-2)=17*-1*2=-34

OpenStudy (anonymous):

owe well thx